# Understanding Optimal Feature Transfer via a Fine-Grained Bias-Variance Analysis

### v1

@@ 5.1 Aligned, Spectrum-Only Case and Connection to PCR
We now turn to understanding how it is possible for the EEP to outperform the OFP. Understanding the solution to Eq. (15) is challenging in general. To build intuition, we first restrict to optimizing only the eigenvalues of $\widehat{\mathbf{B}}$ and fixing its eigenvectors to align with those of $\mathbf{\Sigma}$. Recall Eq. (4) and denote the eigen-decompositions by
@@
for orthonormal eigenbases $\{\mathbf{u}_{i}\}_{i=1}^{p}$ and $\{\hat{\mathbf{q}}_{i}\}_{i=1}^{p}$ (rows of $\widehat{\mathbf{Q}}$). We restrict to minimizing $\mathfrak{R}^{\mathsf{avg}}$ over the eigenvalues $\{\hat{d}_{i}^{2}\}_{i=1}^{p}$ and regularization parameters $\bm{\lambda}$ while fixing the eigenvectors
@@
Recall from Eq. (8), the interpretation of $\hat{\bm{\Gamma}}^{-1/2}=\sum_{i=1}^{p}\hat{r}^{-1/2}_{i}\mathbf{u}_{i}\mathbf{u}_{i}^{\top}$ as a linear featurization on the data matrix $\mathbf{X}$ before regressing on $\mathbf{y}$. This has a natural connection to classical principal component regression (PCR), which projects to the top-$k$ eigenvectors of $\mathbf{\Sigma}$ ($k<n$), commonly referred to as principal components (PCs), via the transform $\hat{\bm{\Gamma}}^{-1/2}=\sum_{i=1}^{k}\mathbf{u}_{i}\mathbf{u}_{i}^{\top}$. Therefore, the optimization of $\hat{r}_{i}$ may be understood as a generalization of PCR with soft feature selection: larger values of $\hat{r}_{i}$ deemphasize the corresponding PC $\mathbf{u}_{i}$, with $\hat{r}_{i}=\infty$ signifying that $\mathbf{u}_{i}$ is not selected. Indeed, choosing $\hat{r}_{i}=1,\forall i\leq k$ and $\hat{r}_{i}=\infty,\forall i>k$ recovers classical PCR.
Observe from Eq. (15) that $\mathfrak{R}^{\mathsf{avg}}$ is a simple function of the variance-component $\mathcal{V}$ and the bias-component $\mathfrak{B}^{\mathsf{avg}}$. We thus first study the properties of $\mathcal{V}$ and $\mathfrak{B}^{\mathsf{avg}}$ with respect to $\{\hat{d}_{i}^{2}\}_{i=1}^{p}$ and $\bm{\lambda}$. The solutions of the separated optimization problems $\min_{\{\hat{r}_{i}\}_{i=1}^{p}}\mathcal{V}\quad\text{and}\quad\min_{\{\hat{r}_{i}\}_{i=1}^{p}}\mathfrak{B}^{\mathsf{avg}}$ can be characterized explicitly, using the results Propositions E.1-E.2 derived in the appendix, and a careful analysis of the Karush–Kuhn–Tucker (KKT) conditions.
@@
We prove the existence and uniqueness of $h_{0}$ and the following theorem in Appendix F.
@@
Suppose $h>n$ (sample-deficient regime) and Eq. (17) holds. Then, $\mathcal{V}$ is minimized by $\hat{r}_{i}=c\eta_{i}$ for $i\leq h$ and any finite $c>0$, to the optimal value $(h/n-1)^{-1}$. Meanwhile, the minimum of $\mathfrak{B}^{\mathsf{avg}}$ undergoes a phase transition w.r.t. $h_{1}$:
@@
Choosing $\hat{\mathbf{q}}_{i}\leftarrow\mathbf{u}_{i},\forall i$ as in Eq. (17) and $\hat{r}_{i}\leftarrow c\eta_{i},\forall i\leq h$ as in Theorem 5.1 is in fact optimal for minimizing $\mathcal{V}$, without a priori restrictions $\hat{\mathbf{q}}_{i}\leftarrow\mathbf{u}_{i},\forall i$. See Proposition F.2 for details.
@@
We defer the algorithmic question of minimizing $\mathfrak{R}^{\mathsf{avg}}$ w.r.t. $\quantity{\hat{r}_{i}}_{i=1}^{p}$ to Appendix E, where we make further observations on the convexity of $\mathcal{V}$ and $\mathfrak{B}^{\mathsf{avg}}$ and provide novel convex programs for optimizing relaxations of $\mathfrak{R}^{\mathsf{avg}}$ and $\mathfrak{R}^{\mathsf{worst}}$.
Remark 5.4 (Technical Novelty).
Our proof strategy follows the following steps: (i) observing the convexity in the bias-component $\mathfrak{B}^{\mathsf{avg}}$ and the variance-component $\mathcal{V}$, (ii) casting the fixed point equation Equation 13 as an equality constraint, (iii) deriving the optimal solution from the KKT conditions. This approach may be of independent interest for studying optimality question in other statistics or machine learning models where the risk asymptotics involve fixed point equations [63, 6, 19, 74].
Recall that $\mathfrak{B}^{\mathsf{avg}}$ characterizes the bias. We first interpret the minimizer of $\mathfrak{B}^{\mathsf{avg}}$ in the soft-selection regime. Note that for $i\leq h_{0}$ we may write Eq. (21) as
@@
Recall that the total variance is proportional to $\mathcal{A}_{1}$, so Theorem 5.1 shows that to control the variance, we should weight the PCs proportional to $\eta_{i}$ regardless of $\theta_{i}$. Larger $\eta_{i}$ therefore have large $\hat{r}_{i}$ and are deemphasized, which is in contrast to classical PCR.
@@ 5.2 Fully-Optimized Representation
Although intractable analytically, $\mathfrak{R}^{\mathsf{avg}}$ can be optimized using backpropagation (see Appendix G). Many of the observations from Section 5.1 carry over to this case in Figure 3, through a combination of eigenvector alignment and spectrum adjustment. By changing $q$, which plays the role of $h_{1}$ when all of $\widehat{\mathbf{B}}$ is optimized, we can transition between the hard- and soft-selection regimes. Similarly, by adjusting the SNR, we can change the relative contributions of $\mathcal{V}$ and $\mathfrak{B}^{\mathsf{avg}}$ to the risk (see Eq. (15)). When the SNR is high, $\mathfrak{B}^{\mathsf{avg}}$ dominates $\mathcal{A}_{1}$, and vice versa.
In the high-SNR, hard-selection regime $\widehat{\mathbf{B}}$’s top-$q$ eigenvectors align with top eigenvectors of $\mathbf{B}^{\star}{\mathbf{B}^{\star}}^{\top}$ to minimize bias (see Eq. (22)). The remaining eigenvectors have a sharp transition to no alignment with $\mathbf{B}^{\star}{\mathbf{B}^{\star}}^{\top}$ and show an inverse relationship with those of $\mathbf{\Sigma}$, as expected from Theorem 5.1. In the low-SNR, hard-selection regime, the alignment to the bottom eigenvectors of $\mathbf{\Sigma}$ is even more pronounced and dominates the alignment to $\mathbf{B}^{\star}{\mathbf{B}^{\star}}^{\top}$ even for the top eigenvectors. The low-SNR, soft-selection regime is similar to the low-SNR, hard selection regime as the alignment to the bottom eigenvectors of $\mathbf{\Sigma}$ dominates and is unaffected by soft vs hard selection. Finally, in the high-SNR, soft-selection regime alignment to $\mathbf{B}^{\star}{\mathbf{B}^{\star}}^{\top}$ is again dominant.
@@
These results interpolate between optimizing $\mathcal{V}$ and $\mathfrak{B}^{\mathsf{avg}}$; see Section I.4 for $\widehat{\mathbf{B}}$’s behavior when optimizing each individually. For example, the spectrum-gap at $q=50$ in Figure 3, (c) may be attributed to ${\mathfrak{B}}^{\mathsf{avg}}$, which, when optimized alone, yields exactly $q=50$ non-zero eigenvalues (Figure 7, (c)), consistent with hard-selection behavior in the spectrum-only case. In contrast, minimizing $\mathfrak{B}^{\mathsf{avg}}$ in the soft-selection regime yields no spectrum-gap at $q$; instead, the number of non-zero eigenvalues ($\sim$120) is smaller than $q=150$ (Figure 7, (f)), consistent with theoretical prediction in the spectrum-only case (i.e. $\mathrm{rank}(\widehat{\mathbf{B}})\equiv h_{0}\leq h_{1}$).
6 Discussion and Future Directions
Appendix A Explicit Expression of the Predictor
Proof of Proposition 2.4.
where
It then follows from standard OLS formulas that
Note that $\hat{\bm{\Lambda}}$ contains matrix quantities
Appendix B Risk Characterization
@@
We prove Theorem 3.3 in this section. As by-products, we provide bounds on asymptotic quantities such as $b_{0},\mathcal{V},\mathfrak{B}$ that will also be useful in later sections; we also provide characterization of semi-classical bias-variance decomposition, which is not covered in the main text.
B.1 Useful Bounds on Asymptotic Quantities
@@
Using Lemma B.1, we prove another lemma that bounds the size of $b_{0}$ defined in (13) and $\mathcal{V},\mathfrak{B}$ defined in (14).
Lemma B.2.
@@
B.2 Characterize Semi-Classical Bias-Variance Decomposition
Before we prove Theorem 3.3, we first characterize the total risk and its semi-classical bias-variance decomposition. The following proposition relies on the main results from [26].
Proposition B.3.
@@ B.2 Characterize Semi-Classical Bias-Variance Decomposition
Proof of Proposition B.3.
Using (8), we have that
@@
where we have used (11) and Lemma B.1. This and other assumptions in 3.1 ensures that [26], Assumption 1 holds and we may directly apply [26], Theorem 2 to characterize $B_{\mathsf{SC}}$ and $V_{\mathsf{SC}}$ above. We conclude the proof by noting that that
@@
B.3 Characterize Fine-grained Bias-Variance Decomposition
Below, we characterize fine-grained bias $B$ and variance components $V_{\mathbf{X}},V_{{\bm{\varepsilon}}},V_{\mathbf{X},{\bm{\varepsilon}}}$, leveraging Proposition B.3 and results from [39], which also proves Theorem 3.3 along with Proposition B.3.
@@ B.3 Characterize Fine-grained Bias-Variance Decomposition
$B$ $=\mathbb{E}_{\mathbf{x}_{\mathrm{new}}}\quantity({\mathbf{x}_{\mathrm{new}}}^{\top}\mathbb{E}_{\mathbf{X},{\bm{\varepsilon}}}\quantity({\hat{\bm{\beta}}}-{\bm{\beta}^{\star}}))^{2}$
@@
where we have used (35) for the last equality.
@@
Note that using (11) and Lemma B.1, we have that
@@
This bound follows from (42), (47) along with Lemma B.1 and B.2, and the fact that $\hat{\mathbf{P}}_{h,\perp}$ is an orthogonal projection as defined in (41).
@@
It remains to prove characterization result for $V_{\mathbf{X}}$ and $V_{\mathbf{X},{\bm{\varepsilon}}}$. It follows from (36), Lemma B.1 and Lemma B.2 that
@@
where $\sigma_{i}(\cdot)$ denotes the $i$-th largest eigenvalue. Here, we used at $(i)$ Von Neumann’s trace inequality, at (ii) the variational representation of eigenvalues and the fact that $\sigma_{\mathrm{min}}(\hat{\mathbf{T}}_{h})=\hat{t}_{\min}^{+}$, at (iii) moment property of inverted Wishart distribution (see e.g. [40], Theorem 2.4.14.) and at (iv) the assumption that $1+M^{-1}<h/n<M$ along with Lemma B.1. Using these bounds, an approach analogous to (40) yields that
@@
Appendix C Downstream Risk Concentration
Lemma C.1 (Hanson-Wright Inequality).
Let $\bm{\xi}$ be some random vector satisfying conditions in Lemma C.2 and $\mathbf{A}$ be any $n\times n$ real-valued matrix. There exists an absolute constant $c>0$ such that for any $x>0$, with probability larger than $1-\exp(-x)$, we have
where $\norm{\cdot}_{F}$ denotes Frobenius norm.
We first prove the following lemma using Hanson-Wright inequality.
Lemma C.2.
Proof of Lemma C.2.
Let us introduce the notation
Then, we can write
[…]
Using the elementary inequality $x/(1+xb_{0})\leq(4b_{0})^{-1}$, Lemma B.1, Lemma B.2 and 3.1, we obtain that
It then follows from 3.1 and the above that
we have that
The required statement follows if we set $x\leftarrow D\cdot\log q$. ∎
We are now ready to prove Proposition 4.7.
Proof of Proposition 4.7.
Using triangle inequality, we have that
[…]
Using the above, Lemma B.1 and Lemma B.2, we obtain
[…]
The result then follows from Theorem 3.3, Lemma C.2 and an application of the union bound.
Appendix D Minimax Optimality
where, with $\sigma_{\mathrm{max}}(\cdot)$ denoting the top eigenvalue,
Instead of tackling the bi-level optimization on the LHS, we may minimize RHS using backpropagation routine developed in Appendix G. For the latter, note that it is standard to differentiate through eigendecomposition and thus $\sigma_{\mathrm{max}}(\cdot)$. See Appendix G for more detailed discussion and pointers to relevant reference. Furthermore, as we will show in Appendix E, $\mathfrak{B}^{\mathsf{worst}}$ may be written into a convex objective under linear constraint under the “spectrum-only” case discussed in Section 5.1; this allows us to design efficient convex programs for convex relaxation of $\mathfrak{R}^{\mathsf{worst}}$.
Definition D.1 (Minimax-optimal pretraining).
The ensuing result provides an justification for Definition D.1, showing that the objective $\mathfrak{R}^{\mathsf{worst}}$, minimized to determine $\widehat{\mathbf{B}}$, tends to approximate the actual worst-case risk.
Proposition D.2.
Proof of Proposition D.2.
Note that
from Theorem 3.3 implies the following
Appendix E Convexity of Spectrum-Only Optimization and Convex Programs
$\hat{\mathbf{q}}_{i}=\mathbf{u}_{i},\forall i\in\{1,\ldots,p\}.$ (57)
which are defined in Section 4 and Appendix D respectively. In this section, we cast the above optimization problems in a more general form
where $f$ is arbitrary function. We show in Section E.1 below that to solve the above, we may solve a linearly constraint optimization problem
where $\tilde{\mathcal{V}},\tilde{\mathfrak{B}}^{\mathsf{avg}}$ and $\tilde{\mathfrak{B}}^{\mathsf{worst}}$ are objectives after applying certain change of variables to $\mathcal{V},\mathfrak{B}^{\mathsf{avg}}$ and $\mathfrak{B}^{\mathsf{avg}}$. We give their definition in Section E.1. A convenient fact that we prove in Section E.2 is that $\tilde{\mathcal{V}},\tilde{\mathfrak{B}}^{\mathsf{avg}}$ and $\tilde{\mathfrak{B}}^{\mathsf{worst}}$ are in in fact convex functions of $\mathbf{x}_{H}$. The above results are important for proving Theorem 5.1. As a byproduct, we also provide convex relaxation of the objectives $\mathfrak{R}^{\mathsf{avg}}$ and
E.1 Equivalence
Given $x_{i}\in[0,1]$ for $i\in\{1,\ldots,p\}$, we define a new objective
where $\mathbf{x}_{H}\in[0,1]^{h}$ with entries $x_{i},i\in H$ and
Proposition E.1 (Equivalence).
Given (17), the solution $\{r_{i}\}_{i=1}^{p}$ of the optimization problem
Proof of Proposition E.1.
Given (17), we have that
where $b_{0}$ is defined as the unique solution of the following equation
$\tilde{\mathcal{F}}\leq\mathcal{F}.$ (65)
E.2 Convexity
Proposition E.2 (Convexity).
Proof of Proposition E.2.
We first note that
$\quantity{\mathbf{x}_{H}\in\mathbb{R}^{h}:h^{-1}\sum_{i\in H}x_{i}=1-n/h}.$
E.3 Algorithmic Implications
The results from Section E.1 and Section E.2 may be leveraged to construct efficient convex algorithms. Using Section E.1 results, we see that to optimize
we only need to solve the following optimization problems
We showed in Section E.2 that $\tilde{\mathcal{V}},\tilde{\mathfrak{B}}^{\mathsf{avg}}$ and $\tilde{\mathfrak{B}}^{\mathsf{worst}}$ are convex by themselves. However, due the presence of the interaction terms $\tilde{\mathfrak{B}}^{\mathsf{avg}}\cdot\tilde{\mathcal{V}}$ and $\tilde{\mathfrak{B}}^{\mathsf{worst}}\cdot\tilde{\mathcal{V}}$, it is uncertain that the objective will remain convex. That said, the objective is in a simple enough form where most constrained local opitmization solvers may be applied. For example, we found that it is quite efficient to solve (66) with the sequential quadratic programming (SQP) routine implemented in Python SciPy package [67].
Therefore, we may consider the following optimization problems
It is easy to verify using Proposition E.2 that the optimization problems (67) are convex program, which can be solved efficiently (see [10] for a review). The solution, obtained from $\hat{r}_{i}\leftarrow c\eta_{i}\quantity(\frac{x_{i}}{1-x_{i}}),i\in H$ will minimize the relaxed upper bounds, $\mathfrak{R}^{\mathsf{avg}}_{\mathsf{rel}}$ and $\mathfrak{R}^{\mathsf{worst}}_{\mathsf{rel}}$, of the original objectives, $\mathfrak{R}^{\mathsf{avg}}$ and $\mathfrak{R}^{\mathsf{worst}}$.
Appendix F Characterize Solutions of Spectrum-Only Optimization
We prove Theorem 3.3 in this section, using Proposition E.2 and Proposition E.1 proved in the previous section. The proof draws extensive use of the Karush-Kuhn-Tuker (KKT) optimality conditions; see [10], Section 5 for a review.

### v2

@@ 5 Structures of the Optimal Representation
In the previous section, we introduced the end-to-end predictor (EEP). This predictor uses a feature matrix $\widehat{\mathbf{B}}=\widehat{\mathbf{B}}^{\mathrm{opt}}$ and regularization parameters $\bm{\lambda}=\bm{\lambda}^{\mathrm{opt}}$ chosen to minimize the asymptotic downstream risk averaged over the downstream ensemble $\mathfrak{R}^{\mathsf{avg}}(\widehat{\mathbf{B}},\bm{\lambda},\mathbf{\Sigma},\mathbf{B}^{\star})$. In this section, we explore the structure of the optimal feature matrix and its connection to the bias-variance decomposition..
@@ 5.1 Aligned, Spectrum-Only Case and Connection to PCR
Understanding the solution to Eq. (12) is challenging in general. To build intuition, we first restrict to optimizing only the eigenvalues of $\widehat{\mathbf{B}}$ and fixing its eigenvectors to align with those of $\mathbf{\Sigma}$. Recall Eq. (4) and denote the eigen-decompositions by
@@
for orthonormal eigenbases $\{\mathbf{u}_{i}\}_{i=1}^{p}$ and $\{\hat{\mathbf{q}}_{i}\}_{i=1}^{p}$ (rows of $\widehat{\mathbf{Q}}$). For the discussion below, we restrict to the spectrum-only case where we minimize $\mathfrak{R}^{\mathsf{avg}}$ over the eigenvalues $\{\hat{d}_{i}^{2}\}_{i=1}^{p}$ and regularization parameters $\bm{\lambda}$ while fixing the eigenvectors
@@
Note that the predictor $\hat{y}$ in Definition 2.1 may be written as $\hat{y}(\mathbf{x}_{\mathrm{new}})=\tilde{\mathbf{x}}_{\mathrm{new}}^{\top}\quantity({{\tilde{\mathbf{X}}}}{{\tilde{\mathbf{X}}}})^{+}{{\tilde{\mathbf{X}}}^{\top}}\mathbf{y},$ where $\tilde{\mathbf{x}}_{\mathrm{new}}:={\mathbf{x}_{\mathrm{new}}^{\top}}\hat{\bm{\Gamma}}^{-1/2}$ and $\tilde{\mathbf{X}}:=\mathbf{X}\hat{\bm{\Gamma}}^{-1/2}$. The matrix quantity $\hat{\bm{\Gamma}}^{-1/2}=\sum_{i=1}^{p}\hat{r}^{-1/2}_{i}\mathbf{u}_{i}\mathbf{u}_{i}^{\top}$ can then be interpreted as a linear featurization on the data matrix $\mathbf{X}$ before regressing on $\mathbf{y}$. This has a natural connection to classical principal component regression (PCR), which projects to the top-$k$ eigenvectors of $\mathbf{\Sigma}$ ($k<n$), commonly referred to as principal components (PCs), via the transform $\hat{\bm{\Gamma}}^{-1/2}=\sum_{i=1}^{k}\mathbf{u}_{i}\mathbf{u}_{i}^{\top}$. Therefore, the optimization of $\hat{r}_{i}$ may be understood as a generalization of PCR with soft feature selection: larger values of $\hat{r}_{i}$ deemphasize the corresponding PC $\mathbf{u}_{i}$, with $\hat{r}_{i}=\infty$ signifying that $\mathbf{u}_{i}$ is not selected. Indeed, choosing $\hat{r}_{i}=1,\forall i\leq k$ and $\hat{r}_{i}=\infty,\forall i>k$ recovers classical PCR.
Observe from Eq. (12) that $\mathfrak{R}^{\mathsf{avg}}$ is a simple function of the variance-component $\mathcal{V}$ and the bias-component $\mathfrak{B}^{\mathsf{avg}}$. We thus first study the properties of $\mathcal{V}$ and $\mathfrak{B}^{\mathsf{avg}}$ with respect to $\{\hat{d}_{i}^{2}\}_{i=1}^{p}$ and $\bm{\lambda}$. The solutions of the separated optimization problems $\min_{\{\hat{r}_{i}\}_{i=1}^{p}}\mathcal{V}\quad\text{and}\quad\min_{\{\hat{r}_{i}\}_{i=1}^{p}}\mathfrak{B}^{\mathsf{avg}}$ can be characterized explicitly, using the results Propositions D.1-D.2 derived in the appendix, and a careful analysis of the Karush–Kuhn–Tucker (KKT) conditions.
@@
We prove the existence and uniqueness of $h_{0}$ and the following theorem in Section 6.2.
@@
Suppose $h>n$ and Eq. (14) holds. Then, $\mathcal{V}$ is minimized by $\hat{r}_{i}=c\eta_{i}$ for $i\leq h$ and any finite $c>0$, to the optimal value $(h/n-1)^{-1}$. Meanwhile, the minimum of $\mathfrak{B}^{\mathsf{avg}}$ undergoes a phase transition w.r.t. $h_{1}$:
@@
Choosing $\hat{\mathbf{q}}_{i}\leftarrow\mathbf{u}_{i},\forall i$ as in Eq. (14) and $\hat{r}_{i}\leftarrow c\eta_{i},\forall i\leq h$ as in Theorem 5.1 is in fact optimal for minimizing $\mathcal{V}$, without a priori restrictions $\hat{\mathbf{q}}_{i}\leftarrow\mathbf{u}_{i},\forall i$. See Proposition 6.5 for details.
@@
We defer the algorithmic question of minimizing $\mathfrak{R}^{\mathsf{avg}}$ w.r.t. $\quantity{\hat{r}_{i}}_{i=1}^{p}$ to Appendix D, where we make further observations on the convexity of $\mathcal{V}$ and $\mathfrak{B}^{\mathsf{avg}}$ and provide novel convex programs for optimizing relaxations of $\mathfrak{R}^{\mathsf{avg}}$ and $\mathfrak{R}^{\mathsf{worst}}$.
We first interpret the minimizer of $\mathfrak{B}^{\mathsf{avg}}$. Firstly, we note that the solution $\{\hat{r}_{i}\}$ is in fact sparse in both the soft-selection regime and strong-selection regime even in absence of sparsity-inducing priors or penalties. In the soft-selection regime, note that for $i\leq h_{0}$ we may write Eq. (18) as
@@
Recall that the total variance is proportional to $\mathcal{V}$, so Theorem 5.1 shows that to control the variance, we should weight the PCs proportional to $\eta_{i}$ regardless of $\theta_{i}$. Larger $\eta_{i}$ therefore have large $\hat{r}_{i}$ and are deemphasized, which is in contrast to classical PCR.
@@ 5.2 Fully-Optimized Representation
Now we would like to lift the assumption (14) of aligned eigenvectors. Although intractable analytically in this case, $\mathfrak{R}^{\mathsf{avg}}$ can be optimized using backpropagation (see Appendix E). Many of the observations from Section 5.1 carry over to this case in Figure 3, through a combination of eigenvector alignment and spectrum adjustment. By changing $q$, we can transition between the hard- and soft-selection regimes. Similarly, by adjusting the SNR, we can change the relative contributions of $\mathcal{V}$ and $\mathfrak{B}^{\mathsf{avg}}$ to the risk (see Eq. (12)). When the SNR is high, $\mathfrak{B}^{\mathsf{avg}}$ dominates $\mathcal{V}$, and vice versa.
In the high-SNR regime, the bias component $\mathfrak{B}^{\mathsf{avg}}$ dominates. Here, $\widehat{\mathbf{B}}$’s top-$q$ eigenvectors align with the leading eigenvectors of $\mathbf{B}^{\star}{\mathbf{B}^{\star}}^{\top}$ to minimize bias, corresponding to the feature selection phenomenon predicted in Theorem 5.1. At the same time, there is a pronounced alignment with the bottom eigenvectors of $\mathbf{\Sigma}$ beyond the $q$th eigenvector, reflecting the impact of the $\mathcal{V}$ component. In the hard selection regime (with $q=50$), a clear spectral gap is present at $q=50$, along with a sharp transition in the alignment pattern before and after the $q$th eigenvector. In contrast, the soft selection regime (with $q=150$) does not have a spectral gap and transition in eigenvector alignment pattern is less sharp. This behavior aligns with the hard and soft-selection phase transition predicted in Theorem 5.1. In the low-SNR regime, the $\mathcal{V}$ component becomes dominant. As a result, we observe a reverse alignment pattern with eigenvectors of $\mathbf{\Sigma}$, as expected from Theorem 5.1.
@@
The above results interpolate between optimizing $\mathcal{V}$ and $\mathfrak{B}^{\mathsf{avg}}$. We now investigate $\widehat{\mathbf{B}}$’s behavior when optimizing each individually. We reproduce Figure 3 under exact settings except that we optimize for $\mathfrak{B}^{\mathsf{avg}}$ only. This gives us Figure 4, (a)-(f). We do the same but optimize for $\mathcal{V}$ only. This gives us Figure 4, (g)-(l). The spectrum-gap at $q=50$ in Figure 3, (c) may be attributed to ${\mathfrak{B}}^{\mathsf{avg}}$, which, when optimized alone, yields exactly $q=50$ non-zero eigenvalues (Figure 4, (c)), consistent with hard-selection behavior in the spectrum-only case. In contrast, minimizing $\mathfrak{B}^{\mathsf{avg}}$ in the soft-selection regime yields no spectrum-gap at $q$; instead, the number of non-zero eigenvalues ($\sim$120) is smaller than $q=150$ (Figure 4, (f)), consistent with theoretical prediction in the spectrum-only case (i.e. $\mathrm{rank}(\widehat{\mathbf{B}})\equiv h_{0}\leq h_{1}$). Observe in (h), (k), the eigenvectors of $\widehat{\mathbf{B}}\widehat{\mathbf{B}}^{\top}$ align with eigenvectors of $\mathbf{\Sigma}$ in exactly the reverse order.
6 Proofs of Main Results
Before presenting the proofs, we briefly outline the key technical contributions of this paper. First, we use random matrix theory to derive asymptotic characterization of fine-grained bias and variance. We first note that that the fine-grained bias-variance components relate to the classical decomposition as follows:
$\mathbb{E}_{\mathbf{X}}B_{\mathsf{SC}}=B+V_{\mathbf{X}},\qquad\mathbb{E}_{\mathbf{X}}V_{\mathsf{SC}}=V_{\mathbf{X},{\bm{\varepsilon}}}+V_{{\bm{\varepsilon}}}.$ (21)
The asymptotic expressions for $\mathbb{E}_{\mathbf{X}}B_{\mathsf{SC}}$ and $\mathbb{E}_{\mathbf{X}}V_{\mathsf{SC}}$ follow from the analysis in [23], and the term $V_{{\bm{\varepsilon}}}$ vanishes due to symmetry. Thus, it suffices to derive the limiting expression for fine-grained bias $B$. We observed that $B$ involves two independent copies of a random matrix quantity rather than a single instance and applied resolvent methods separately to each copy (see (37)).
The second major contribution lies in the optimality analysis of asymptotic risk in the aligned case. A key insight is that the limiting expressions for fine-grained bias and variance can be reformulated as a convex optimization problem over the spectrum of the feature matrix, where the fixed point equation (10) emerging from RMT is cast as a linear constraint. This reduces the analysis to solving the KKT conditions of the convex problem. The primary challenge is in conjecturing a solution that satisfies these conditions—particularly when optimizing $\widehat{\mathbf{B}}$, where the solution exhibits sparsity and undergoes a phase transition. Our approach may be of independent interest for studying optimality question in other statistics or machine learning models where the risk asymptotics involve fixed point equations [59, 6, 18, 70].
6.1 Proof of Theorem 3.3
@@ 6 Proofs of Main Results
We prove Theorem 3.3 in this section. As by-products, we provide bounds on asymptotic quantities such as $b_{0},\mathcal{V},\mathfrak{B}$ that will also be useful in later sections; we also provide characterization of fine-grained bias-variance decomposition.
6.1.1 Preliminary Bounds on Asymptotic Quantities
@@
Using Lemma 6.1, we prove another lemma that bounds the size of $b_{0}$ defined in (10) and $\mathcal{V},\mathfrak{B}$ defined in (11).
Lemma 6.2.
@@
6.1.2 Characterization of Fine-Grained Bias-Variance Decomposition
Before we prove Theorem 3.3, we first characterize the total risk and its fine-grained bias-variance decomposition. The following proposition relies on the main results from [23].
Proposition 6.3.
@@ 6.1.2 Characterization of Fine-Grained Bias-Variance Decomposition
Proof of Proposition 6.3.
Note that have that
@@
where we have used (8) and Lemma 6.1. This and other assumptions in 3.1 ensures that [23], Assumption 1 holds and we may directly apply [23], Theorem 2 to characterize $B_{\mathsf{SC}}$ and $V_{\mathsf{SC}}$ above. We conclude the proof by noting that that
@@
6.1.3 Characterization of Fine-grained Bias-Variance Decomposition
Below, we characterize fine-grained bias $B$ and variance components $V_{\mathbf{X}},V_{{\bm{\varepsilon}}},V_{\mathbf{X},{\bm{\varepsilon}}}$, leveraging Proposition 6.3 and results from [37], which also proves Theorem 3.3 along with Proposition 6.3.
@@ 6.1.3 Characterization of Fine-grained Bias-Variance Decomposition
$B$ $=\mathbb{E}_{\mathbf{x}_{\mathrm{new}}}\quantity({\mathbf{x}_{\mathrm{new}}}^{\top}\mathbb{E}_{\mathbf{X},{\bm{\varepsilon}}}\quantity({\hat{\bm{\beta}}}-{\bm{\beta}^{\star}}))^{2}$ (36)
@@
where we have used (30) for the last equality. Note that the key difference between fine-grained bias (36) and classical bias $B_{\mathsf{SC}}$ defined in (31) is that $B$ is the expectation over
${{\bm{\beta}^{\star}}}^{\top}\hat{\bm{\Gamma}}^{1/2}\hat{\mathbf{W}}_{h}^{\top}(\mathbf{I}-\hat{\mathbf{P}}_{h}^{(1)})\hat{\mathbf{T}}_{h}(\mathbf{I}-\hat{\mathbf{P}}_{h}^{(2)})\hat{\mathbf{W}}\hat{\bm{\Gamma}}^{1/2}{{\bm{\beta}^{\star}}}$ (37)
for two independent copies $\hat{\mathbf{P}}_{h}^{(1)},\hat{\mathbf{P}}_{h}^{(2)}$ of random matrix $\hat{\mathbf{P}}_{h}$. Our strategy is then to apply resolvent method twice for the two independent copies.
@@
Note that using (8) and Lemma 6.1, we have that
@@
This bound follows from (39), (44) along with Lemma 6.1 and 6.2, and the fact that $\hat{\mathbf{P}}_{h,\perp}$ is an orthogonal projection as defined in (38).
@@
It remains to prove characterization result for $V_{\mathbf{X}}$ and $V_{\mathbf{X},{\bm{\varepsilon}}}$. It follows from (31), Lemma 6.1 and Lemma 6.2 that
@@
where $\sigma_{i}(\cdot)$ denotes the $i$-th largest eigenvalue. Here, we used at $(i)$ Von Neumann’s trace inequality, at (ii) the variational representation of eigenvalues and the fact that $\sigma_{\mathrm{min}}(\hat{\mathbf{T}}_{h})=\hat{t}_{\min}^{+}$, at (iii) moment property of inverted Wishart distribution (see e.g. [38], Theorem 2.4.14.) and at (iv) the assumption that $1+M^{-1}<h/n<M$ along with Lemma 6.1. Using these bounds, an approach analogous to (35) yields that
@@
6.2 Proof of Theorem 5.1
We prove Theorem 5.1 in this section, using Proposition D.2 and Proposition D.1 proved in the previous section. The proof draws extensive use of the Karush-Kuhn-Tucker (KKT) optimality conditions; see [10], Section 5 for a review.